OpenStudy (anonymous):

I am trying to solve the equation (5x-4)(4x+8)=0 using factoring. I keep getting stuck on this problem. Could someone please answer and explain how you got your answer? Thanks you! :)

jimthompson5910 (jim_thompson5910):

If A*B = 0, then either A = 0 or B = 0

jimthompson5910 (jim_thompson5910):

So using that idea, we can go from (5x-4)(4x+8)=0 to 5x-4 = 0 or 4x+8 = 0

jimthompson5910 (jim_thompson5910):

Do you see how to finish up?

OpenStudy (anonymous):

Yes, thank you

jimthompson5910 (jim_thompson5910):

no problem

OpenStudy (decentnabeel):

\[\mathrm{Using\:the\:Zero\:Factor\:Principle:}\] \[\mathrm{Solve\:}\:5x-4=0:\quad x=\frac{4}{5}\] \[\mathrm{Solve\:}\:4x+8=0:\quad x=-2\] \[\mathrm{The\:final\:solutions\:\to\:the\:quadratic\:equation\:are:}\] \[x=\frac{4}{5},\:x=-2\] the answer is .. \[\left(5x-4\right)\left(4x+8\right)=0\quad :\quad x=\frac{4}{5},\:x=-2\]